OpenStudy (anonymous):

graph both of the following lines on the same graph plane. Name each line. state the slope and any intercepts of each equation A) 2x-3=7 B) -15y=45

OpenStudy (anonymous):

Before you graph these, it is probably a good first step to solve for the one variable in each equation. So, for (A), solve for x so you have "x = ..." and for (B), "y = ......" Then you can graph more easily.

OpenStudy (anonymous):

i have a problem with my answers not sure it they are right. M=0, B=-3, xint= 3/2

OpenStudy (anonymous):

That doesn't look right... Let's go through it. 2x - 3 = 7 2x - 3 + 3 = 7 + 3 <<-- add 3 to both sides 2x = 10 2x/2 = 10/2 <<-- divide both sides by 2 x = 5

OpenStudy (anonymous):

seems weird, right? Where's the normal y = mx + b form? This line is a special case... it's a vertical line, and slope is undefined.

OpenStudy (anonymous):

was i right for B atleast

OpenStudy (anonymous):

The vertical line is x = 5, so it cuts the x-axis at x = 5 and goes up and down vertically from there. The problem asks for slope and "any intercepts", so the slope is "undefined", and you answer the intercepts part by saying "the x-axis intercept, or x-intercept, is 5". If you are used to regular y = mx + b lines, this question is almost a trick question. It's important to learn it, but the way they asked it was a little tricky :(

OpenStudy (anonymous):

yea i know

OpenStudy (anonymous):

For question B, again, solve for the variable first... -15y = 45 -15y/(-15) = 45/(-15) y = -3 So this one may look a little more like a y = mx + b line... y = mx + b y = - 3 Notice there's no "x" term... that's because slope is actually 0. When you have a line equation with no "x" term, and just "y = (some number)", it is a horizontal line (slope = 0), and it runs through the y-intercept of whatever that number was.... in this case, -3.

OpenStudy (anonymous):

So, keep in mind... 3 types of line equations: x = 5 <<--- vertical line, slope undefined, x axis intercept at 5 y = 5 <<--- horizontal line, slope = 0, y-axis intercept at 5 y = 2x + 5 <<---- sloped line with slope = 2, y axis intercept at 5